OpenStudy (anonymous):

A farmer feeds his cows 200 pounds of feed each day and has 700 pounds of feed in his barn. Another farmer feeds his cow 350 pounds of feed each day and has 1000 pounds of feed in his barn. a.in how many days will the two farmers have the same amount of feed left? b.does your answer make sense?explain. c. How would your answer change if both farmers got an additional 1000 pounds of feed.

OpenStudy (anonymous):

Set the problem up like this:\[700-200x=1000-350x \implies 150x=300 \implies x=2\]In two days, both farmers will have four hundred pounds of feed.

OpenStudy (anonymous):

I think it makes perfect sense.

OpenStudy (anonymous):

The adjusted problem for 1000 lb more at each farm is similar:\[1700-200x=2000-350x\]Amazingly (or maybe not.....) the answer is still two days, although they will each have 1400 lb instead of 400 lb.

OpenStudy (anonymous):

can you show me how to do it then/